For the first week of April, each evening we recorded the number of coins in the fountain. And we have the number of coins is linear function of how many days of April has passed. But when we went out to celebrate our discovery we loss most of our data in a freak ketchup accident. Here's all that left.

April Coins in the fountain
1
2. 229
3
4
5
6. 297
7


question: what is the m value for this function
1
Expert's answer
2020-04-27T19:11:06-0400

m (slope) ="\dfrac{f(x_1)-f(x_2)}{x_1-x_2}" ="\dfrac{297-229}{6-2}=\dfrac{68}{4}=17"